Assume that the time a student stays in School is normally distributed with a mean of 5 hours and a standard deviation of 0.5 hours. Every day, Ian stays in School for 5.5 hours. What proportion of students stays less than 5.5 hours?
1
Expert's answer
2021-09-27T16:14:49-0400

"\mu = 5 \\

\sigma=0.5 \\

P(X<5.5) = P(Z< \frac{5.5-5.0}{0.5}) \\

= P(Z< 1) \\

= 0.8413"

(Using Z distribution table)